satellite73 (satellite73):

A wagon train that is one mile long travels one mile at a constant rate. During the same time period, a horse rides at a constant rate from the front of the wagon train to the rear, and then back to the front. How far did the horse ride?

OpenStudy (zarkon):

do you have a solution or are you actually looking for a solution?

OpenStudy (anonymous):

i spent an hour on it this morning, and yes i have a solution you can google it and get one, but honestly i did not understand the yahoo answer, so i redid it differently thought it was kind of cute because it made me think

OpenStudy (zarkon):

I just found the time it takes to get to the back and then the time it takes to get back to the front. then \[d=v_1\cdot t_1+v_2\cdot t_2\]

OpenStudy (anonymous):

ok maybe half an hour, but still

OpenStudy (zarkon):

though \[v_1=v_2=h\] h=speed of the horse

OpenStudy (beginnersmind):

I think I got the logic, but can't do the algebra :(

OpenStudy (zarkon):

well if you have an answer...tell me if I got it right... \[\frac{2h^2}{(w+h)(h-w)}\] where h = speed of horse and w= speed of wagon

OpenStudy (anonymous):

the answer is an actual number

OpenStudy (beginnersmind):

I think you can get a specific answer without knowing the actual speed

OpenStudy (zarkon):

that seems weird to me...unless I am not understanding the question

OpenStudy (foolaroundmath):

the answer should be 3 miles. Looking for a way to explain that other than intuition

OpenStudy (anonymous):

no it is not three, although i did get three the first time

OpenStudy (zarkon):

what if the horse moves at the same rate as the wagon?

OpenStudy (foolaroundmath):

oops ... its 2miles .. i'll draw a figure.. hopefully that'll help

OpenStudy (zarkon):

does the wagon stop after 1 mile?

OpenStudy (anonymous):

don't forget that the horse travels front to back and back to front at the same time the train goes one mile

OpenStudy (anonymous):

so it cannot travel at the same rate as the train

OpenStudy (zarkon):

oh ... ok...well that is different

OpenStudy (beginnersmind):

If w is the speed of the wagon and h is the speed of the horse it takes t=1/w time for the wagon to travel one mile. The horse takes 1/(h+w) time to get from the start to the end and 1(h-w) to get from the end to the front. So we get 1/h = 1/(h+w)+1/(h/w). This should be solved for h/w

OpenStudy (anonymous):

in fact that was the key to the solution, i picked 1 for the speed of the train, then solved for the speed of te horse

OpenStudy (foolaroundmath):

 The train moves 1 mile. The initial position of the train (left rectangle) and initial position of the horse (rear of the train = extreme left) are shown. After the train has moved 1 mile, the train's new position is drawn (The right rectangle). and the final position of the horse is shown (front of the train = extreme right). Now its easy to see that the horse traveled 1+1 = 2miles